7(3-x)+1=5x-2
21-7x+1=5x-2
24=12x
24/12=2
12x/12=x
x=2

Answer by samgagner18(1)

7(3-x)+1=5x-2
21-7x+1=5x-2
24=12x
24/12=2
12x/12=x
x=2